A bus travels at 40kmph for a distance of 80 km and at 60kmph for a distance of 180km. The average speed is __
1
Expert's answer
2011-07-05T07:46:00-0400
Average speed = total distance / total time

T1 = 80 [km]/40 [kmph] = 2 [h]
T2 = 180 [km] / 60 [kmph] = 3 [h]
Total time = T1 + T2 = 2 + 3 = 5 [h]
Total distance = 80 + 180 = 260 [km]

Average speed = 260 [km] / 5 [h] = 52 [kmph]